Problem: Coffee Brand A =8.00 per pound. Brand B =5.00 per pound. The two brands mixed together = 6.00 per pound. How many pounds of Brand A will be needed to make a blend of a total of 50 pounds?
Answer by mananth(16946)

---------price ---------------- quantity
Brand A 8 ---------------- x pounds
Brand B 5 ------ 50 - x pounds
Mixture 6.00% ---------------- 50

8 x + 5 ( 50 - x ) = 6 * 50

8x+250-5x=300
8x-5x=300-250
3x=50
/3
x=16.67 pounds Brand A
33.33 pounds Brand B